8.14 Comparing Investment Criteria Wii Brothers, a game manufacturer, has a new idea for an adventure game. It can market the game either as a traditional board game or as an interactive DVD, but not both. Consider the following cash flows of the two mutually exclusive projects for the company. Assume the discount rate for both projects is 10 percent. Main Page Year Board Game DVD 0 $(950.00) $(2,100.00) 1 $700.00 $1,500.00 2 $550.00 $1,050.00 3 $130.00 $450.00 Discount Rate Payback NPV IRR "a. Based on the payback period rule, which project should be chosen? b. Based on the NPV, which project should be chosen? c. Based on the IRR, which project should be chosen?"
Answer #1

1.
Board Game
Payback=1+(950-700)/550=1.454545455

DVD
Payback=1+(2100-1500)/1050=1.571428571

Choose Board Game

2.
Board Game
NPV=-950+700/1.1+550/1.1^2+130/1.1^3=238.580015

DVD
NPV=-2100+1500/1.1+1050/1.1^2+450/1.1^3=469.4966191

Choose DVD

3.
Board Game
IRR:
0=-950+700/(1+IRR)+550/(1+IRR)^2+130/(1+IRR)^3
IRR=27.5065%

DVD
IRR:
0=-2100+1500/(1+IRR)+1050/(1+IRR)^2+450/(1+IRR)^3
IRR=25.0928%

Choose Board Game